# CLEAR: Calibrated Learning for Epistemic and Aleatoric Risk

### v1

@@ Abstract
Accurate uncertainty quantification is critical for reliable predictive modeling, especially in regression tasks. Existing methods typically address either aleatoric uncertainty from measurement noise or epistemic uncertainty from limited data, but not necessarily both in a balanced way. We propose CLEAR, a calibration method with two distinct parameters, $\gamma_{1}$ and $\gamma_{2}$, to combine the two uncertainty components for improved conditional coverage. CLEAR is compatible with any pair of aleatoric and epistemic estimators; we show how it can be used with (i) quantile regression for aleatoric uncertainty and (ii) ensembles drawn from the Predictability–Computability–Stability (PCS) framework for epistemic uncertainty. Across 17 diverse real-world datasets, CLEAR achieves an average improvement of 28.2% and 17.4% in the interval width compared to the two individually calibrated baselines while maintaining nominal coverage. This improvement can be particularly evident in scenarios dominated by either high epistemic or high aleatoric uncertainty. Code: https://github.com/Unco3892/clear
@@ 1 Introduction
Uncertainty quantification is essential for building reliable machine learning systems [1, 2]. Despite their impressive capabilities, modern machine learning methods can give a false sense of reliability; therefore, producing valid prediction intervals remains an open problem. Calibration [3] and conformal methods [4, 5, 6, 7, 8, 9] adjust prediction intervals to obtain marginal coverage (that is, covering a certain percentage of the data on average). However, they may suffer from poor conditional coverage, meaning well-calibrated coverage at the individual or subgroup level [10]. In particular, under distribution shift or model misspecification, conditional coverage can degrade substantially, especially in extrapolation regions. Most conformal methods, such as conformalized quantile regression (CQR) [9], only capture aleatoric uncertainty while ignoring epistemic uncertainty.
@@
The combination of epistemic and aleatoric uncertainty is not new: Bayesian methods have long incorporated both components [17, 18], and several recent approaches have also explored this decomposition in the context of conformal prediction [19, 20, 21]. However, to the best of our knowledge, existing methods either do not explicitly distinguish between the two types of uncertainty or implicitly fix the combination ratio, for instance, setting $\lambda=1$ as in [22, 17, 18], or fixing $\gamma_{1}=1$ as in [19]. This fixed choice may be suboptimal, as the relative importance of each uncertainty type varies with the data distribution and prediction task (see Appendix A for more details on related work).
@@ 1.1 Contributions
    We are the first to introduce two calibration parameters $\gamma_{1}$ and $\gamma_{2}$ to balance the scale of epistemic and aleatoric uncertainty on the validation dataset.
@@
    We are the first to combine the ensemble perturbation intervals via PCS framework ([23] and [24, Chapter 13]) with the CQR aleatoric uncertainty estimator, and empirically show the strengths of this combination.
@@
    We take a step towards large-scale benchmarking of uncertainty quantification by extending the PCS-UQ framework of [23] to methods based on quantile regression.
The remainder of the paper is organized as follows: Section 2 introduces the CLEAR methodology. Section 3 outlines the experimental setup, including synthetic simulations, real-world datasets, baselines, and metrics. Section 4 presents results and a case study of the PCS pipeline. Section 5 concludes with limitations and future directions.
@@ 2.1 Problem scenario
where $\alpha\in(0,1)$ is for example $\alpha=0.05$. In order to construct $C$, data $\mathcal{D}=\{(X_{i},Y_{i}),i=1,\dots,n\}$ can be split into train and calibration subsets $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{cal}}$. On the training data, a first estimate of $C$ can be constructed, and then we can use data from $\mathcal{D}_{\text{cal}}$ to calibrate $C$ such that (2) is satisfied.
@@ 2.2 Epistemic uncertainty
The traditional machine learning approach trains a predictive algorithm on a single version of the cleaned/preprocessed dataset and uses the best-performing algorithm (compared using the validation set) for future predictions. While theoretically sound in the infinite-sample limit, this approach ignores the uncertainty stemming from finite sample size and model choice (epistemic uncertainty). Various methods have been proposed to estimate this uncertainty, including Deep Ensembles [22], MC dropout in NN [27], Orthonormal Certificates [13], NOMU [28], BNNs [29], Laplace Approximation [30], to name a few.
Estimating Epistemic Uncertainty via PCS: In practice, additional sources of uncertainty arise from subjective choices in data cleaning, imputation, and dataset construction, which we also consider as extended epistemic uncertainty. The Predictability, Computability, and Stability (PCS) framework [24, 23, 31] offers a holistic point of view on the data-science-life-cycle, though it does not explicitly model aleatoric uncertainty. Following [24, Chapter 13], we obtain an ensemble of $m$ estimators $\hat{f}_{1},\dots,\hat{f}_{m}$ as follows:
@@
    We split the data into $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{val}}$.
@@
    We create $N_{1}$ differently preprocessed versions of the data and define $N_{2}$ different models (such as linear regression, random forest, neural networks, etc.). We train models for all $N_{1}\times N_{2}$ combinations on $\mathcal{D}_{\text{train}}$ and pick the top-$k$ based on their performance on $\mathcal{D}_{\text{val}}$.
@@
    We refit each of the top-$k$ models on $b$ bootstrap samples $\mathcal{D}_{\text{train}}^{1},\dots,\mathcal{D}_{\text{train}}^{b}$ of $\mathcal{D}_{\text{train}}$ to obtain an ensemble of $m=k\times b$ estimators $\hat{f}_{1},\dots,\hat{f}_{m}$.
Taking the point-wise median of $\hat{f}_{1},\dots,\hat{f}_{m}$, yields the final PCS estimate $\hat{f}$, and the point-wise $\alpha/2$ and $21-\alpha/$ 2 quantiles (denoted $\hat{f}_{\alpha/2}$ and $\hat{f}_{1-\alpha/2}$, respectively) define the (uncalibrated) uncertainty band. The widths of this interval, denoted as $\hat{q}^{\text{epi}}_{1-\alpha/2}(x):=\hat{f}_{1-\alpha/2}(x)-\hat{f}(x)$ and $\hat{q}^{\text{epi}}_{\alpha/2}(x):=\hat{f}(x)-\hat{f}_{\alpha/2}(x)$, quantifies the pre-calibrated epistemic uncertainty.
For implementation details of PCS uncertainty quantification, see Section C.2. A recent extension of PCS uncertainty quantification [23] uses the combined data set $\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{val}}$ for training and calibrates uncertainty based on out-of-bag data. We expect that CLEAR could benefit from incorporating such refinements. Although most of our experiments only cover the modeling step of the PCS framework, as in [23], in Section 4.3 we also account for uncertainty due to data cleaning and pre-processing on a case study following Chapter 13 in [24].
@@ 2.3 Aleatoric uncertainty
Aleatoric uncertainty can be estimated by modeling the conditional distribution of the outcome given the inputs. Common approaches include direct conditional quantile regression [32], either parametric or nonparametric, such as smooth quantile regression [33] and quantile random forests (QRF) [34]; heteroskedastic models that estimate input-dependent noise levels $\sigma(x)$ under Gaussian assumptions [35]; and distributional regression techniques such as simultaneous quantile regression [13]. More flexible alternatives include conditional density estimation and deep generative models such as conditional generative adversarial networks, conditional variational autoencoders, or diffusion models [36, 37, 38].
In this work, we estimate aleatoric uncertainty using quantile regression models $\hat{r}_{\alpha/2},\hat{r}_{0.5},\hat{r}_{1-\alpha/2}$ (selected in step 2 from PCS) trained on the residuals $Y_{i}-\hat{f}(X_{i})$. This approach offers improved stability: underfitting quantile regression directly on the $y$-values can severely distort aleatoric uncertainty estimates. In contrast, extreme underfitting on residuals, at worst, corresponds to assuming homoscedastic noise, which can be an acceptable bias. This improves the stability with respect to hyperparameters. To further improve the stability, we apply a PCS-inspired bagging strategy. From the aggregated estimators of $\hat{r}_{\alpha},\hat{r}_{0.5},\hat{r}_{1-\alpha}$, we compute $\hat{q}^{\text{ale}}_{1-\alpha/2}(x):=\hat{r}_{1-\alpha/2}(x)-\hat{r}_{0.5}(x)$ and $\hat{q}^{\text{ale}}_{\alpha/2}(x):=\hat{r}_{0.5}(x)-\hat{r}_{\alpha/2}(x)$. Overall, it is computationally efficient and straightforward to implement.
@@ 2.4 CLEAR: combining aleatoric & epistemic uncertainty
Given a fixed ratio $\lambda=\frac{\gamma_{2}}{\gamma_{1}}$, we compute $\gamma_{1}$ on a held-out calibration set using the standard split conformal prediction procedure. While the natural choice $\gamma_{1}=\gamma_{2}$ may seem appealing, it is often suboptimal. This is because the relative contribution of aleatoric and epistemic uncertainty can vary across datasets, and the corresponding estimators may differ substantially in scale and precision.
To choose $\lambda$ from data, we evaluate a grid of positive values $\Lambda$. For each candidate $\lambda\in\Lambda$, we construct the (calibrated) interval $C_{\lambda}$. To ensure the best trade-off between uncertainty sources, we select $\lambda^{\star}$ such that $C_{\lambda^{\star}}$ performs best under the chosen metric on $\mathcal{D}_{\text{val}}$. We have chosen quantile loss [32] (defined in Algorithm 1, also known as pinball loss) as a simple metric to balance both coverage and width. However, any other metric can also be used. As a proper scoring rule, quantile loss incentivizes truthfulness from a theoretical perspective. This procedure is summarized in Algorithm 1.
In the case that either the estimation of epistemic or the estimation of aleatoric uncertainty terribly fails, $\lambda^{\star}$ can compensate this failure to some extent by giving more weight on the other estimator. In the other (hopefully more likely) case, that both estimators $\hat{q}^{\text{epi}}$ and $\hat{q}^{\text{ale}}$ perform reasonably well (up to incorrect scaling), $\lambda^{\star}$ can provide valuable insights beyond predictive inference: A very large value $\lambda^{\star}\frac{\hat{q}_{1-\alpha/2}^{\text{epi}}(x)+\hat{q}_{\alpha/2}^{\text{epi}}(x)}{\hat{q}_{1-\alpha/2}^{\text{ale}}(x)+\hat{q}_{\alpha/2}^{\text{ale}}(x)}\gg 1$ hints that epistemic uncertainty dominates the uncertainty for $x$, and is expected to reduce if you collect more training observations close to $x$ (or by making stronger and better modeling assumptions). Vice versa, a small value $\lambda^{\star}\frac{\hat{q}_{1-\alpha/2}^{\text{epi}}(x)+\hat{q}_{\alpha/2}^{\text{epi}}(x)}{\hat{q}_{1-\alpha/2}^{\text{ale}}(x)+\hat{q}_{\alpha/2}^{\text{ale}}(x)}\ll 1$ hints that aleatoric uncertainty dominates for $x$ which cannot be reduced by collecting more training observations. In some cases, the aleatoric uncertainty can be reduced by changing the problem formulation by increasing the dimension of $x$ (i.e., by measuring more co-variates for the training and validation dataset and for the new datapoint $x$).
@@
contains at least $\lceil(1-\alpha)(|\mathcal{D}_{\text{cal}}|+1)\rceil$ of points in $\mathcal{D}_{\text{cal}}$. See Appendix B for a detailed implementation.
@@
9: For any metric in Section C.3 (e.g., quantile loss), evaluate $C_{\lambda}(x)$ on $\mathcal{D}_{\text{val}}$ and set
@@
where $\text{QuantileLoss}(\mathcal{D}_{\text{val}},C_{\lambda}):=\sum_{i\in\mathcal{D}_{\text{val}}}\left[QL_{\alpha/2}\big{(}Y_{i},l(X_{i})\big{)}+QL_{1-\alpha/2}\big{(}Y_{i},u(X_{i})\big{)}\right]/2$, with $l(x),u(x)$ denoting the bounds of $C_{\lambda}(x)$, and $QL_{\tau}(y,q)=(y-q)\big{(}\tau-\mathds{1}_{(-\infty,q]}(y)\big{)}$.
@@
The consistency of the predictors implies $\hat{q}_{\tau}^{\text{epi}}(x)\to 0$, and the consistency of the quantile estimators implies $\hat{q}_{\tau}^{\text{ale}}(x)\to q_{\tau}^{\text{ale}}(x)$ point-wise. Hence, the estimated pre-calibrated intervals converge to their population analogues. Since $\sup_{\lambda\in\Lambda}\lambda\hat{q}_{\tau}^{\text{epi}}(x)\to 0$ converges to 00 uniformly over $\lambda\in\Lambda$ (due to compactness of $\Lambda$), therefore $C(x)$ is asymptotically equal to true conditional quantiles $[{r}_{\alpha/2}(x),{r}_{1-\alpha/2}(x)]$; and necessarily $\lim_{|\mathcal{D}_{\text{cal}}|\to\infty}\gamma_{1}=1$, as is the case in classical CQR (see e.g. [8, Section 5]). ∎
Note that these assumptions are satisfied by the finite grid $\Lambda$ used in our implementation and by many base models, including tree-based methods and neural networks. We elaborate on the theoretical implications in Appendix G, where formal guarantees for marginal coverage are provided in Lemma G.2.

### v2

@@ Abstract
Accurate uncertainty quantification is critical for reliable predictive modeling. Existing methods typically address either aleatoric uncertainty due to measurement noise or epistemic uncertainty resulting from limited data, but not both in a balanced manner. We propose CLEAR, a calibration method with two distinct parameters, $\gamma_{1}$ and $\gamma_{2}$, to combine the two uncertainty components and improve the conditional coverage of predictive intervals for regression tasks. CLEAR is compatible with any pair of aleatoric and epistemic estimators; we show how it can be used with (i) quantile regression for aleatoric uncertainty and (ii) ensembles drawn from the Predictability–Computability–Stability (PCS) framework for epistemic uncertainty. Across 17 diverse real-world datasets, CLEAR achieves an average improvement of 28.2% and 17.4% in the interval width compared to the two individually calibrated baselines while maintaining nominal coverage. Similar improvements are observed when applying CLEAR to Deep Ensembles (epistemic) and Simultaneous Quantile Regression (aleatoric). The benefits are especially evident in scenarios dominated by high aleatoric or epistemic uncertainty. [^1] Code: https://github.com/Unco3892/clear
@@ 1 Introduction
Uncertainty quantification (UQ) is essential for building reliable machine learning systems [1, 2]. Despite their impressive capabilities, modern machine learning methods can give a false sense of reliability; therefore, producing valid prediction intervals remains an open problem. Calibration [3] and conformal methods [4, 5, 6, 7, 8] adjust prediction intervals to obtain marginal coverage (that is, covering a certain percentage of the data on average). However, they may suffer from poor conditional coverage, meaning well-calibrated coverage at the individual or subgroup level [9]. In particular, under distribution shift or model misspecification, conditional coverage can degrade substantially, especially in extrapolation regions. Most conformal methods, such as conformalized quantile regression (CQR) [10], only capture aleatoric uncertainty while ignoring epistemic uncertainty.
@@
The combination of epistemic and aleatoric uncertainty is not new: Bayesian methods have long incorporated both components [17, 18], and several recent approaches have also explored this decomposition in the context of conformal prediction [19, 20, 21]. However, to the best of our knowledge, existing methods either do not explicitly distinguish between the two types of uncertainty or implicitly fix the combination ratio, for instance, setting $\lambda=1$ [22, 17, 18], or fixing $\gamma_{1}=1$ [19]. This fixed choice may be suboptimal, as the relative importance of each uncertainty type varies with the data distribution and prediction task (see Appendix A for more details on related work).
@@ 1.1 Contributions
    We are the first to introduce two calibration parameters $\gamma_{1}$ and $\gamma_{2}$, to balance the scales of epistemic and aleatoric uncertainty on the validation dataset.
@@
    We are the first to combine the ensemble perturbation intervals of the PCS framework [23] with the CQR aleatoric uncertainty estimator, and empirically show the strengths of this combination.
@@
    We conduct large-scale UQ benchmarking for several models on 17 regression datasets.
@@ 2.1 Problem scenario
where $\alpha\in(0,1)$ is for instance $\alpha=0.05$. In order to construct $C$, data $\mathcal{D}=\{(X_{i},Y_{i}),i=1,\dots,n\}$ can be split into train and calibration subsets $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{cal}}$. On the training data, a first estimate of $C$ can be constructed, and then we can use data from $\mathcal{D}_{\text{cal}}$ to calibrate $C$ such that (2) is satisfied.
@@ 2.2 Epistemic uncertainty
The traditional machine learning approach trains a predictive algorithm on a single version of the cleaned/preprocessed dataset and uses the best-performing algorithm (compared using the validation set) for future predictions. While theoretically sound in the infinite-sample limit, this approach ignores the uncertainty stemming from finite sample size and model choice (epistemic uncertainty). Various methods have been proposed to estimate this uncertainty, including Deep Ensembles [22], MC dropout in NN [26], Orthonormal Certificates [13], NOMU [27], BNNs [28] and Laplace Approximation [29], among others.
Estimating Epistemic Uncertainty via PCS: In practice, additional sources of uncertainty arise from subjective choices in data cleaning, imputation, and dataset construction, which we also consider as extended epistemic uncertainty. The Predictability, Computability, and Stability (PCS) framework [23] offers a holistic point of view on the data-science-life-cycle, without explicitly modeling aleatoric uncertainty. One can obtain an ensemble of $m$ estimators $\hat{f}_{1},\dots,\hat{f}_{m}$ as follows:
@@
    Split the data into $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{val}}$.
@@
    Create $N_{1}$ differently preprocessed versions of the data and define $N_{2}$ different models (e.g., linear regression, random forest, neural networks). Then, train models for all $N_{1}\times N_{2}$ combinations on $\mathcal{D}_{\text{train}}$ and pick the top-$k$ based on their performance on $\mathcal{D}_{\text{val}}$.
@@
    Refit each of the top-$k$ models on $b$ bootstrap samples $\mathcal{D}_{\text{train}}^{1},\dots,\mathcal{D}_{\text{train}}^{b}$ of $\mathcal{D}_{\text{train}}$ to obtain an ensemble of $m=k\times b$ estimators $\hat{f}_{1},\dots,\hat{f}_{m}$.
Taking the point-wise median of $\hat{f}_{1},\dots,\hat{f}_{m}$, yields the final PCS estimate $\hat{f}$, and the point-wise $\alpha/2$ and $1-\alpha/2$ quantiles (denoted $\hat{f}_{\alpha/2}$ and $\hat{f}_{1-\alpha/2}$, respectively) define the (uncalibrated) uncertainty band. The widths of this interval, denoted as $\hat{q}^{\text{epi}}_{1-\alpha/2}(x):=\hat{f}_{1-\alpha/2}(x)-\hat{f}(x)$ and $\hat{q}^{\text{epi}}_{\alpha/2}(x):=\hat{f}(x)-\hat{f}_{\alpha/2}(x)$, quantifies the pre-calibrated epistemic uncertainty. [30] extends this by using the combined data set $\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{val}}$ for training and calibrating uncertainty based on out-of-bag data. The focus is only on the modeling step of the framework. We later demonstrate—through a case study in Section 4.3—that accounting for uncertainty from data cleaning and pre-processing is also important, and how the uncertainty contributions can be adapted and improved using our approach.
@@ 2.3 Aleatoric uncertainty
Aleatoric uncertainty can be estimated by modeling the conditional distribution of the outcome given the inputs. Common approaches include direct conditional quantile regression [31], either parametric or nonparametric, such as smooth quantile regression [32] and quantile random forests (QRF) [33]; heteroskedastic models that estimate input-dependent noise levels $\sigma(x)$ under Gaussian assumptions [34]; and distributional regression techniques such as simultaneous quantile regression [13]. More flexible alternatives include conditional density estimation and deep generative models such as conditional generative adversarial networks [35], conditional variational autoencoders [36] and diffusion models [37].
In this work, we estimate aleatoric uncertainty using quantile regression models $\hat{r}_{\alpha/2},\hat{r}_{0.5},\hat{r}_{1-\alpha/2}$ (selected in Line 2 from PCS) trained on the residuals $Y_{i}-\hat{f}(X_{i})$. This approach offers improved stability: underfitting quantile regression directly on the $y$-values can severely distort aleatoric uncertainty estimates. In contrast, extreme underfitting on residuals, at worst, corresponds to assuming homoskedastic noise, which can be an acceptable bias. This improves the stability with respect to hyperparameters. To further improve the stability, we apply a PCS-inspired bagging strategy, by taking the empirical median
$\hat{q}^{\text{ale}}_{1-\alpha/2}(x):=\text{Median}\left[\left(\hat{r}_{1-\alpha/2}(x)-\hat{r}_{0.5}(x)\right)^{+}\right]\text{ and }\hat{q}^{\text{ale}}_{\alpha/2}(x):=\text{Median}\left[\left(\hat{r}_{0.5}(x)-\hat{r}_{\alpha/2}(x)\right)^{+}\right]$
over the ensemble members. Overall, it is computationally efficient and straightforward to implement.
@@ 2.4 CLEAR: combining aleatoric & epistemic uncertainty
Given a fixed ratio $\lambda=\frac{\gamma_{2}}{\gamma_{1}}$, we compute $\gamma_{1}$ on a held-out calibration set using the standard split conformal prediction procedure. While the natural choice $\gamma_{1}=\gamma_{2}$ may seem appealing, it is often suboptimal. The relative contribution of aleatoric and epistemic uncertainty can vary across datasets, and the corresponding estimators may differ substantially in scale and precision when $\gamma_{1}=\gamma_{2}$.
To choose $\lambda$ from data, we evaluate a grid of positive values $\Lambda$. For each candidate $\lambda\in\Lambda$, we construct the (calibrated) interval $C_{\lambda}$. To ensure the best trade-off between uncertainty sources, we select $\lambda^{\star}$ such that $C_{\lambda^{\star}}$ performs best under the chosen metric on $\mathcal{D}_{\text{val}}$. We have chosen quantile loss [31] (defined in Algorithm 1, also known as pinball loss) as a simple metric to balance both coverage and width. However, any other metric can also be used. As a proper scoring rule, quantile loss incentivizes truthfulness from a theoretical perspective (see Section B.3). This procedure is summarized in Algorithm 1.
Parameter $\lambda^{\star}$ balances aleatoric and epistemic uncertainties: if one estimator fails, $\lambda^{\star}$ compensates by re-weighting the other. When both estimators $\hat{q}^{\text{epi}}$ and $\hat{q}^{\text{ale}}$ are reliable (up to scaling), $\lambda^{\star}$ is interpretable. A large ratio
$\lambda^{\star}\tfrac{\hat{q}_{1-\alpha/2}^{\text{epi}}(x)+\hat{q}_{\alpha/2}^{\text{epi}}(x)}{\hat{q}_{1-\alpha/2}^{\text{ale}}(x)+\hat{q}_{\alpha/2}^{\text{ale}}(x)}\gg 1$
indicates that epistemic uncertainty dominates at $x$ (reducible with more data or stronger assumptions), while a small ratio $\ll 1$ indicates aleatoric uncertainty dominates (irreducible, though sometimes mitigated by adding covariates).
@@
contains at least $\lceil(1-\alpha)(|\mathcal{D}_{\text{cal}}|+1)\rceil$ of points in $\mathcal{D}_{\text{cal}}$. See Section B.5 for implementation.
@@
9: For any metric in Section D.5 (e.g., quantile loss), evaluate $C_{\lambda}(x)$ on $\mathcal{D}_{\text{val}}$ and set
@@
where $\text{QuantileLoss}(\mathcal{D}_{\text{val}},C_{\lambda}):=\frac{1}{|\mathcal{D}_{\text{val}}|}\sum_{i\in\mathcal{D}_{\text{val}}}\left[QL_{\alpha/2}\big(Y_{i},l(X_{i})\big)+QL_{1-\alpha/2}\big(Y_{i},u(X_{i})\big)\right]/2$, with $l(x),u(x)$ denoting the bounds of $C_{\lambda}(x)$, and $QL_{\tau}(y,q)=(y-q)\big(\tau-\mathds{1}_{(-\infty,q]}(y)\big)$.
@@
The consistency of the predictors implies $\hat{q}_{\tau}^{\text{epi}}(x)\to 0$, and the consistency of the quantile estimators implies $\hat{q}_{\tau}^{\text{ale}}(x)\to q_{\tau}^{\text{ale}}(x)$ point-wise. Hence, the estimated pre-calibrated intervals converge to their population analogues. Since $\sup_{\lambda\in\Lambda}\lambda\hat{q}_{\tau}^{\text{epi}}(x)\to 0$ converges to 0 uniformly over $\lambda\in\Lambda$ (due to compactness of $\Lambda$), therefore $C(x)$ is asymptotically equal to true conditional quantiles ${r}_{\alpha/2}(x),{r}_{1-\alpha/2}(x)]$; and necessarily $\lim_{|\mathcal{D}_{\text{cal}}|\to\infty}\gamma_{1}=1$, as is the case in classical CQR (see e.g. [8, Section 5]). ∎
Note that these assumptions are satisfied by the finite grid $\Lambda$ used in our implementation and by many base models, including tree-based methods and neural networks. We elaborate on the theoretical implications in Appendix B, with formal guarantees for marginal coverage in Lemma B.2.